OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. 3x^2 + 5x + 1 = 0

OpenStudy (tkhunny):

Can you identify the traditional definitions of a, b, and c?

OpenStudy (anonymous):

Im not sure I understand what you mean..

OpenStudy (studygurl14):

The Quadratic Formula: Where... \[ax^2+bx+c=0\] ... \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (studygurl14):

Take the values of a, b, and c from the equation, and plug into the quadratic formula, and solve for x

OpenStudy (anonymous):

no not you @StudyGurl14 I was talking to @tkhunny

OpenStudy (studygurl14):

If... \[3x^2+5x+1=0\] ...and... \[ax^2+bx+c=0\] ...what are the values of a, b, and c?

OpenStudy (anonymous):

idk how to figure it out

OpenStudy (studygurl14):

Look. I'll give you a little bit more help, but that's it. You have to figure out the rest on your own. \[3x^2=ax^2\] So a= ?

OpenStudy (tkhunny):

You will have to do better than that. If you cannot do this identification and substitution, after having it laid out in front of you, then you are in the wrong class. Look at the two equations given by StudyGurl14 and identify a, b, and c as they have been described. It is time for you to show some work. We really cannot help you if you really have no idea whatsoever.